Air Travel. A turbo-jet flies 50 mph faster than a superprop plane. If a turbo-jet goes 2000 mi in 3 hr less time than it takes the super-prop to go 2800 mi, find the speed of each plane.
Answer by mananth(16946)

let the speed of superprop plane be x mph
the turbo speed will be x+50 mph
..
distance traveled by superprop = 2800
time taken by superprop = 2800/x
..
the turbo travels a distance of 2000 miles
speed = x+50 mph
time = 2000/x+50
Time taken by superprop - time taken by turbo = 3 hours
2000/x-2800/(x+50)= 3
LCM = x(x+50)
2000(x+50)-2800x/ x(x+50)=3
2000x+100000-2800x=3x(x+50)
-800x +100000= 3x^2+150x
3x^2+950x-100000=0
find the roots using quadratic formula
x1=83.33mph speed of superprop
x2=-400 ignore
speed of turbo = 83.33 + 50= 133.33 mph